OpenStudy (anonymous):

please solve? I give medals. 4b/9 - b/5=11, solve for b. a.40 b.45 c.50 d.55

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large \frac{4b}{9} - \frac{b}{5} = 11\] \[\Large \frac{5*4b}{5*9} - \frac{9*b}{9*5} = 11\] \[\Large \frac{20b}{45} - \frac{9b}{45} = 11\] \[\Large \frac{20b-9b}{45} = 11\] \[\Large \frac{11b}{45} = 11\] and I'll let you finish up

OpenStudy (anonymous):

its 45, i used my fancy calculator :P

jimthompson5910 (jim_thompson5910):

very good, here's how you do it without a calculator \[\Large \frac{4b}{9} - \frac{b}{5} = 11\] \[\Large \frac{5*4b}{5*9} - \frac{9*b}{9*5} = 11\] \[\Large \frac{20b}{45} - \frac{9b}{45} = 11\] \[\Large \frac{20b-9b}{45} = 11\] \[\Large \frac{11b}{45} = 11\] \[\Large 11b = 11*45\] \[\Large b = \frac{11*45}{11}\] \[\Large b = \frac{\cancel{11}*45}{\cancel{11}}\] \[\Large b = 45\]